# PhraseVAE and PhraseLDM: Latent Diffusion for Full-Song Multitrack Symbolic Music Generation

### v1

@@ Abstract
This technical report presents a new paradigm for full-song symbolic music generation. Existing symbolic models operate on note-attribute tokens and suffer from extremely long sequences, limited context length, and weak support for long-range structure. We address these issues by introducing PhraseVAE and PhraseLDM, the first latent diffusion framework designed for multitrack symbolic music. PhraseVAE compresses variable-length polyphonic note sequences into compact 64-dimensional phrase-level representations with high reconstruction fidelity, allowing efficient training and a well-structured latent space. Built on this latent space, PhraseLDM generates an entire multi-track song in a single pass without any autoregressive components. The system eliminates bar-wise sequential modeling, supports up to 128 bars of music (8 minutes in 64 bpm), and produces complete songs with coherent local texture, idiomatic instrument patterns, and clear global structure. With only 45M parameters, our framework generates a full song within seconds while maintaining competitive musical quality and generation diversity. Together, these results show that phrase-level latent diffusion provides an effective and scalable solution to long-sequence modeling in symbolic music generation. We hope this work encourages future symbolic music research to move beyond note-attribute tokens and to consider phrase-level units as a more effective and musically meaningful modeling target. [^1] Demos and code: https://www.oulongshen.xyz/midi_ldm.
@@ 1 Introduction
Despite the centrality of songs in musical practice, most music generation systems still operate at the segment level, producing short excerpts rather than complete musical works wang2024whole. This disconnect arises not from a lack of demand, but from the inherent difficulty of modeling long-form symbolic music. A full song typically spans tens or hundreds of bars, involves multiple sections with distinct structural functions, and expresses musical ideas through long-range development rather than isolated moments. Consequently, generating a song requires capturing not only local note-to-note coherence but also global structure, such as sectional form, thematic recurrence, harmonic pacing, texture arrangement, and long-term melodic evolution. Bridging this gap between short-form generative models and real-world musical artifacts motivates the need for more effective full-song symbolic music generation.
Modeling music is an important topic in generative AI, yet full-song symbolic music generation remains relatively underexplored; most existing work instead focuses on segment-level generation. Existing approaches typically rely on sub-note-level tokenization, which we call note-attribute token in the paper—where a single note event is decomposed into multiple tokens—and autoregressive decoding, such as huang2020pop; von2023figaro; ouunifying; wang2025notagen; lu2023musecoco. This design leads to extremely long token sequences, which introduces several inherent limitations. First, the computational cost becomes prohibitive: for instance, generating a 32-bar sample with MuseCoco lu2023musecoco can take several minutes, whereas typical users may expect results within seconds. Second, many systems support only segment-level generation because their supported context length are insufficient to model an entire song ouunifying; zhao2023accomontage; yang2025meteor. Third, the excessive sequence length makes it difficult to learn long-range musical structure, such as maintaining coherent melodic variation across repeated sections, especially under limited training data zhao2023accomontage.
@@ 2.1 Semantics of Music
In symbolic music, this semantic richness does not exist at the note or sub-note level. Most symbolic music models—including nearly all autoregressive ones—encode each note event by decomposing it into multiple note-attribute tokens, where each token represents only a single aspect of the event huang2020pop; qu2024mupt; wang2025notagen; wu2024melodyt5; pasquier2025midi; ouunifying; von2023figaro. For example, in the REMI-z tokenization scheme ouunifying, a token such as o-12 merely indicates that an onset occurs at the first beat of a bar. It provides no information about the pitch, instrument, timbre, duration, expressive intent, or the musical role of the note (e.g., whether it belongs to a melodic line, comping pattern, padding texture, or arpeggio). Ironically, our previous model in ouunifying learns a 768-dimensional embedding for that token whose semantic content is simply “something begins here.” While the efficiency issues of such representations will be discussed later, the key point is that an individual note-attribute token carries almost no semantic meaning on its own, making it a poor unit for modeling music.
@@ 3.2.1 AE with Multi-Query Compression
To ensure reconstruction quality, we start with building an autoencoder (AE) without KL regularization kingma2013auto. Our AE is a sequence-to-sequence model implemented with an encoder–decoder Transformer vaswani2017attention. The input and output of the model are the same sequence, since the model learns to compress and reconstruct it. As in standard encoder–decoder setups, the encoder produces a sequence of hidden states, and the decoder autoregressively reconstructs the sequence during inference. The difference is that the decoder does not attend to the full encoder states. Instead, it receives only a compressed representation of hidden states through a bottleneck.
@@ 3.3.1 Model Structure
There are several implementation details for handling variable-length songs. During training, all latent sequences are padded to the maximum supported length, which is 512 latents (128 bars). This allows the model to generate a fixed-length latent sequence, while different songs may have different actual lengths. To let the model learn where a song should end, we introduce a special end-of-song latent in the VAE training stage, represented by a dedicated token. When training the diffusion model, the latent corresponding to this end-of-song phrase serves as the boundary indicator. If a song is shorter than the maximum length, we pad the remaining positions using this special latent as well. During decoding, the output token sequence from the VAE is truncated at the first occurrence of the end-of-song token. We do not apply attention masks inside the LDM’s self attention. Positional encoding is crucial for indicating the location and ordering of latents, and we adopt rotary position embeddings similar to Stable Audio.
@@ 4.2.3 VAE with Bottleneck
Models with 128-dimensional latents or larger can almost perfectly reconstruct the music. (We will show later that such large latents are also suboptimal due to memorization issues.) A 32-dimensional bottleneck is too restrictive, achieving only 95.1% $\mathbf{F1}_{\mathrm{op}}$. In symbolic music, onset or pitch errors are highly perceptible, often causing reconstructed notes to shift to incorrect timing or incorrect frequencies, which is musically unacceptable. For this reason we did not adopt 32 dimensions. Instead, we choose 64 dimensions, which is compact while still providing high reconstruction fidelity.
@@ 4.3.2 Metrics for Generation Quality
PhraseFID. This metric evaluates phrase-level generation quality. We directly compute FID score heusel2017gans directly in the PhraseVAE latent space between latents of generated songs and that of the training set. It measures (1) how well the model generates phrases that resemble those in the training data, and (2) the phrase-level musical quality, with the assumption that phrases closer to human created phrase have higher quality. A lower PhraseFID indicates that generated phrases follow a distribution closer to real data and are hence better in quality.
@@ 4.3.6 Qualitative Results
Structure-Unconditioned Models Often Produce Coherent Song Structures. Even without structure conditioning, the model frequently generates musically meaningful large-scale structures. For instance, many samples contain clear intros and outros with characteristic sparse textures or cadential gestures, as in Figure 7. Between melodic sections, the model often inserts lead-instrument fills that function as transitions. Although chorus/verse boundaries are not always as distinct as in real songs, we still observe repeated sections that indicate implicit structure planning.
@@ 5.2 Validation Loss
5.3 Latent Dimension
Initially, we attempted to train the LDM using an embedding dimension equal to the VAE latent dimension. Under this configuration, the model was unable to fit even a single-song toy dataset. Increasing the d_model size significantly alleviated this issue, whereas increasing depth alone did not help. For future work, we recommend setting the LDM model dimension to at least twice the latent dimensionality to ensure stable training.

### v2

@@ Abstract
This technical report presents a new paradigm for full-song symbolic music generation. Existing symbolic models operate on note-attribute tokens and suffer from extremely long sequences, limited context length, and weak support for long-range structure. We address these issues by introducing PhraseVAE and PhraseLDM, the first latent diffusion framework designed for full-song multitrack symbolic music. PhraseVAE compresses an arbitrary variable-length polyphonic note sequence into a single compact 64-dimensional phrase-level latent representation with high reconstruction fidelity, allowing a well-structured latent space and efficient generative modeling. Built on this latent space, PhraseLDM generates an entire multi-track song in a single pass without any autoregressive components. The system eliminates bar-wise sequential modeling, supports up to 128 bars of music (8 minutes in 64 bpm), and produces complete songs with coherent local texture, idiomatic instrument patterns, and clear global structure. With only 45M parameters, our framework generates a full song within seconds while maintaining competitive musical quality and generation diversity. Together, these results show that phrase-level latent diffusion provides an effective and scalable solution to long-sequence modeling in symbolic music generation. We hope this work encourages future symbolic music research to move beyond note-attribute tokens and to consider phrase-level units as a more effective and musically meaningful modeling target. [^1] Demos and code: https://www.oulongshen.xyz/midi_ldm.
@@ 1 Introduction
Despite the centrality of songs in musical practice, most music generation systems still operate at the segment level, producing short excerpts rather than complete musical works wang2024whole. This disconnect arises not from a lack of demand, but from the inherent difficulty of modeling long-form symbolic music. A full song typically spans tens or hundreds of bars, involves multiple sections with distinct structural functions, and expresses musical ideas through long-range development rather than isolated moments. Consequently, generating a song requires capturing not only local note-to-note coherence but also global structure, such as sectional form, thematic recurrence, harmonic pacing, texture arrangement, and long-term melodic evolution. Bridging this gap between short-form generative models and real-world musical works motivates the need for more effective full-song symbolic music generation.
Modeling music is an important topic in generative AI, yet full-song symbolic music generation remains relatively underexplored. Existing approaches typically rely on sub-note-level tokenization, which we call note-attribute token in the paper—where a single note event is decomposed into multiple tokens—and autoregressive decoding, such as huang2020pop; von2023figaro; ouunifying; wang2025notagen; lu2023musecoco. This design leads to extremely long token sequences, which introduces several inherent limitations. First, the computational cost becomes prohibitive: for instance, generating a 32-bar sample with MuseCoco lu2023musecoco can take several minutes, whereas typical users may expect results within seconds. Second, many systems support only segment-level generation because their supported context length are insufficient to model an entire song ouunifying; zhao2023accomontage; yang2025meteor. Third, the excessive sequence length makes it difficult to learn long-range musical structure, such as maintaining coherent melodic variation across repeated sections, especially under limited training data zhao2023accomontage.
@@ 2.1 Semantics of Music
In symbolic music, this semantic richness does not exist at the note or sub-note level. Most symbolic music models—including nearly all autoregressive ones—encode each note event by decomposing it into multiple note-attribute tokens, where each token represents only a single aspect of the event huang2020pop; qu2024mupt; wang2025notagen; wu2024melodyt5; pasquier2025midi; ouunifying; von2023figaro. For example, in the REMI-z tokenization scheme ouunifying, a token such as o-12 merely indicates that an onset occurs at the second beat of a bar. It provides no information about the pitch, instrument, timbre, duration, expressive intent, or the musical role of the note (e.g., whether it belongs to a melodic line, comping pattern, padding texture, or arpeggio). Ironically, our previous model in ouunifying learns a 768-dimensional embedding for that token whose semantic content is simply “something begins here.” While the efficiency issues of such representations will be discussed later, the key point is that an individual note-attribute token carries almost no semantic meaning on its own, making it a poor unit for modeling music.
@@ 3.2.1 AE with Multi-Query Compression
To ensure reconstruction quality, we start with building an autoencoder (AE) without KL regularization kingma2013auto. Our AE is a sequence-to-sequence model implemented with an encoder–decoder Transformer vaswani2017attention. The input and output of the model are the same phrase sequence (REMI-z track sequence), since the model learns to compress and reconstruct it. As in standard encoder–decoder setups, the encoder produces a sequence of hidden states, and the decoder autoregressively reconstructs the sequence during inference. The difference is that the decoder does not attend to the full encoder states. Instead, it receives only a compressed representation of hidden states through a bottleneck.
@@ 3.3.1 Model Structure
There are several implementation details for handling variable-length songs. During training, all latent sequences are padded to the maximum supported length, which is 512 latents (128 bars). This allows the model to generate a fixed-length latent sequence, while different songs may have different actual lengths. To let the model learn where a song should end, we introduce a special end-of-song latent in the VAE training stage, represented by a dedicated token. When training the diffusion model, the latent corresponding to this end-of-song phrase serves as the boundary indicator. If a song is shorter than the maximum length, we pad the remaining positions using this special latent as well. During decoding, the output token sequence from the VAE is truncated at the first occurrence of the end-of-song token. We do not apply attention masks inside the LDM’s self attention. Positional encoding is crucial for indicating the location and ordering of latents, and we adopt rotary position embeddings following Stable Audio.
@@ 4.2.3 VAE with Bottleneck
Models with 128-dimensional latents or larger can almost perfectly reconstruct the input music; however, as we will show later, such large latent spaces are suboptimal because they make the subsequent latent diffusion model prone to memorization. In contrast, a 32-dimensional bottleneck is overly restrictive, achieving only 95.1% $\mathbf{F1}_{\mathrm{op}}$. In symbolic music, onset or pitch errors are highly perceptible: even small inaccuracies can cause notes to shift to incorrect timings or pitches, rendering the reconstruction immediately musically unacceptable. For this reason, we do not adopt a 32-dimensional latent space. Instead, we choose a 64-dimensional bottleneck, which remains compact while still providing high reconstruction fidelity.
@@ 4.3.2 Metrics for Generation Quality
PhraseFID. This metric evaluates phrase-level generation quality. We directly compute FID score heusel2017gans in the PhraseVAE latent space between latents of generated songs and that of the training set. It measures (1) how well the model generates phrases that resemble those in the training data, and (2) the phrase-level musical quality, with the assumption that phrases closer to human created phrases have higher quality. A lower PhraseFID indicates that generated phrases follow a distribution closer to real data and are hence better in quality.
@@ 4.3.6 Qualitative Results
Structure-Unconditioned Models Often Produce Coherent Song-Level Structure. Even without structure conditioning, the model frequently generates musically meaningful song-level structure. For instance, many samples contain clear intros and outros with characteristic sparse textures or cadential gestures, as in Figure 7. Between melodic sections, the model often inserts lead-instrument fills that function as transitions. Although chorus/verse boundaries are not always as distinct as in real songs, we still observe repeated sections that indicate implicit structure planning.
@@ 5.1 Beta Value for PhraseVAE
5.2 Latent Space Smoothness and Interpolation Behavior
(a) $\alpha=0.0$
(b) $\alpha=0.3$
(c) $\alpha=0.4$
(d) $\alpha=0.5$
(e) $\alpha=0.6$
(f) $\alpha=0.7$
(g) $\alpha=0.8$
(h) $\alpha=1.0$
Figure 9: Latent space interpolation results of PhraseVAE. Each subfigure shows the decoded phrase obtained by linearly interpolating between two phrase-level latent vectors with interpolation weight $\alpha$.
A desirable property of a variational latent representation is that semantic interpolation in latent space should correspond to smooth and meaningful transformations in the data domain. To examine this property, we perform linear interpolation between two randomly selected phrase-level latent vectors produced by PhraseVAE and decode the interpolated latents back into symbolic music.
Figure 9 visualizes eight interpolation points with interpolation weights ranging from 0.0 to 1.0. The decoded results exhibit clear and consistent smoothness: musical attributes such as pitch, onset timing, duration, and instrument identity change gradually rather than abruptly. At each interpolation step, only a subset of note properties is modified, while the overall rhythmic contour and local textural patterns are largely preserved.
Notably, even when the interpolation traverses regions associated with different instruments, the decoded phrases remain musically coherent. Local note groupings and phrase-level “shapes” are often retained, suggesting that the latent space captures higher-level musical structure rather than surface-level token statistics. This behavior indicates that PhraseVAE learns a continuous and well-structured latent manifold aligned with musically meaningful dimensions.
Such smooth interpolation behavior is particularly important for downstream generative modeling. It implies that neighboring latent points correspond to perceptually similar musical phrases, reducing the risk of abrupt semantic jumps during diffusion-based generation. These results further support that PhraseVAE provides a suitable latent representation for long-range symbolic music modeling.
@@ 5.3 Validation Loss
5.4 LDM Model Dimension
Initially, we attempted to train the LDM using a model dimension equal to the VAE latent dimension. Under this configuration, the model was unable to fit even a single-song toy dataset. Increasing the d_model size significantly alleviated this issue, whereas increasing depth alone did not help. For future work, we recommend setting the LDM model dimension to at least twice the latent dimensionality to ensure stable training.